# Anticipating Gaming to Incentivize Improvement: Guiding Agents in (Fair) Strategic Classification

### v1

@@ 6.2 Effects of fairness interventions
We now proceed with discussing the effect of imposing a fairness constraint $\mathcal{C}$ when selecting the optimal threshold (as formalized in (3)), on the contrast between the unfair policies (characterized in Lemma 4 and 5) with the corresponding fair policies (characterized in Lemmas 6 and 7). Similar to Section 6, we do so when all agents opt for the same response type, and for Types 1 and 3 of agents’ best responses. Our comparison, outlined in the following Corollary, is a direct consequence of the proof of Proposition 8 under the stated assumptions.
@@
    Decreases (increases) manipulation by unqualified advantaged (disadvantaged) agents ($\bm{\Psi}^{0}_{a,(i)}(\hat{\theta}^{\mathcal{C}}_{a})\leq\bm{\Psi}^{0}_{a,(i)}(\hat{\theta}^{U}_{a})$ and $\bm{\Psi}^{0}_{b,(i)}(\hat{\theta}^{\mathcal{C}}_{b})>\bm{\Psi}^{0}_{b,(i)}(\hat{\theta}^{U}_{b}),~{}i\in\{1,3\}$),
@@
In our context, Corollary 9 argues that in its attempt to attain fair outcomes, the firm decreases improvement incentives of the disadvantaged group $b$, while increasing the improvement incentives for the advantaged one; this can be viewed as a negative consequence of fairness interventions on shaping agents’ strategic responses. As we elaborate through numerical experiments in Section 7, this negative by-product is more severe if the firm is non-strategic (i.e., unaware of agents’ strategic responses).
@@ 7 Numerical Experiments
In this section, we will use numerical experiments to support our analytical findings and observations from Section 6. We focus on affordable improvement (Type 1) equilibria which contain a mix of manipulation and improvement responses; similar observations can be made for costly improvement (Type 3) equilibria and are provided in Appendix A.5. We will start with a population consisting of a single demographic group to highlight the main observations from Proposition 8 (including under some relaxations of its assumptions).
We will then presents experiments on a population with two distinct demographic groups, placing emphasis on the impacts of fairness interventions, contrasting the induced strategic behavior in each demographic group, and analyzing the induced post-strategic qualification rates across groups at equilibrium.
Single-group experiment setup.
We start with a population of agents from a single group $s$, with truncated Gaussian feature distributions. In particular, we let $G^{0}_{s}\sim\text{TruncNorm}([20,60],40,15^{2})$ and $G^{1}_{s}=\text{TruncNorm}([53,113],83,15^{2})$. We vary the (pre-strategic) qualification rate $\alpha_{s}$ in $(0,1)$. For the characteristics of the strategic actions, we set $C_{M,s}=0.1$, and $C_{I,s}=0.6$. The boost distributions are truncated Gaussian distributions, parameterized as follows: $\tau^{y}_{M,s}(b)\sim\text{TruncNorm}([10,50],30,22^{2}),$, $\tau^{0}_{I,s}(b)\sim\text{TruncNorm}([37,79],58,15^{2})$, and $\tau^{1}_{I,s}(b)\sim\text{TruncNorm}([40,80],60,15^{2}).$ We draw $n=1000$ total agents from the feature distributions, with the number drawn from in each label determined from the selected $\alpha_{s}$. For each $\alpha_{s}$, we run the experiment $50$ times, and report the average across these runs.
@@
In Figure 3, we explore the impacts of anticipating agents’ strategic behavior (in the absence of any fairness interventions). Figure 3(a) shows that the strategic optimal thresholds are greater than the optimal non-strategic threshold at all levels of $\alpha_{s}$; this is consistent with Proposition 8. Figure 3(b) supports this by showing that a strategic firm’s utility is greater than the non-strategic one’s. Lastly, consistent with parts (ii)-(iv) of Proposition 8, we observe that the anticipation of agents’ strategic behavior reduces the impacts of manipulation by the unqualified agents (Fig. 3(c)) while increasing improvement by them (Fig. 3(d)). Simultaneously, this choice leads to both more manipulation (Fig. 3(e)) and (weakly) more improvement (Fig. 3(f)) by the qualified agents, both of which also benefit the firm. Figure 4 further supports these claims, by showing that post-strategic qualification rates $\hat{\alpha}_{s}$ are higher under the strategic firm’s policy compared to the non-strategic firm’s policy. In other words, strategic firms succeed at incentivizing agents to opt for improvement, with the impacts being more pronounced when the group is initially majority unqualified (low $\alpha_{s}$). Finally, we also note that all these observations hold even though the assumption $\overline{x}^{0}_{s}+(\mathbb{T}^{y}_{M,s})^{-1}(C_{I,s}-C_{M,s})<\mu^{0}_{s}+\underline{b}^{0}_{I,s}$ is not satisfied in the current experiment setup, hinting that Proposition 8 may hold beyond its current assumptions.
(a) Optimal thresholds comparison
(b) Optimal utility comparison
(c) $\bm{\Psi}^{0}_{s,(1)}$ (Manipulation) comparison
(d) $\bm{\Phi}^{0}_{s,(1)}$ (Improvement) comparison
(e) $\bm{\Psi}^{1}_{s,(1)}$ (Manipulation) comparison
(f) $\bm{\Phi}^{1}_{s,(1)}$ (Improvement) comparison
Two-group numerical illustration setup.
Strategic vs. non-strategic fair firms.
The more notable and newly observed distinction is when we compare the impact of the type of fairness intervention (DP vs. TPR) on the difference between strategic and non-strategic firms. In particular, there is a more significant change in DP-fair thresholds when the firm is non-strategic. To see why, consider the DP-Fair policy selected by a non-strategic firm (Fig. 5(a)). Such a firm does not account for agents’ strategic responses, simply lowering the decision threshold on group $b$ and increasing it on group $a$, in order to satisfy the fairness constraint. For group $b$, this DP-Fair policy is in fact so low that it accepts many agents (by default) and accepts others with even lower features who opt for manipulation. That is why $\bm{\Phi}^{0}_{b,(1)}(\theta^{\mathcal{C}}_{b})$ (resp. $\bm{\Psi}^{0}_{b,(1)}(\theta^{\mathcal{C}}_{b})$) decreases (resp. increases) compared to those under the unfair non-strategic policy, as illustrated by the solid red lines in Figures 6 the second and the fourth subplots. In contrast, a strategic firm lowers its fair thresholds on group $b$ much less drastically (Fig. 5(b)), as it realizes that parity between selection rates (as required by DP) can be achieved by a combination of adjusting the thresholds and driving agents’ best-responses. In fact, we can observe that fairness desiderata are achieved with very small changes to the decision thresholds when accounting for agents’ strategic responses, therefore leading to far less loss in utility for the firm.
(a) Non-strategic firm’s utility
(b) Strategic firm’s utility
@@
Impacts on agents’ post-strategic qualification rates.
@@
To further explore the impacts of a strategic policy on agents’ behavior under different fairness interventions (unfair, DP, and TPR), in Figure 7, we contrast the post-strategic $\hat{\alpha}_{s}$ qualification rates for each group. First, as also noted earlier, we note that the strategic firm is more successful at incentivizing improvement, as evidenced by the higher $\hat{\alpha}_{s}$ in both groups $a$ and $b$ compared to the non-strategic firm (in Fig. 7, blue bars are consistently higher than red bars). This increase is even more significant for group $b$, as this group has a lower qualification rate to begin with.
Perhaps more interestingly, we observe that if a firm does not account for strategic responses by agents, fairness interventions can have a negative implication as they decrease improvement by the disadvantaged group $b$ (in Fig. 7, the red bars are lower under the fair policies compared to the unfair policy). This is because fairness interventions typically lower the threshold for group $b$ compared to the unfair policy, which in turn drives the adoption of manipulation over improvement decisions. In comparison, a strategic firm recognizes that fairness goals can be met by a combination of lowering the threshold and incentivizing improvement actions, ultimately leading to higher improvement rates in the disadvantaged group $b$.
@@ 8 Conclusion and Future Work
We proposed a Stackelberg game model to study strategic classification, where a firm deploys a (fair) classifier and agents strategically respond by adjusting their true qualification states and/or observable features to increase their chances of acceptance. We model both different costs and stochastic efficacy for the agents’ manipulation and improvement actions. We find that anticipating strategic behavior can allow the firm to not only curb manipulation behavior (as also noted in prior work) but can also incentivize agents to opt for improvement decisions. Specifically, we find that a strategic firm chooses its policy to incentivize improvement by unqualified agents (driving them to improve both their qualification states and observable features), while still allowing for manipulation by some qualified agents (who do not have sufficiently high observable features to be selected otherwise). We further show that if a firm chooses to impose a fairness constraint on its decision rule, while remaining unaware of agents’ strategic responses, this choice will typically increase (resp. lower) the thresholds for advantaged (resp. disadvantaged) groups, in turn increasing (resp. decreasing) their incentives for improvement. In terms of its impact on a disadvantaged (majority-unqualified) group, this can be viewed as a negative byproduct of fairness interventions. In contrast, a strategic firm understands that desired fairness goals can be attained through a combination of adjusting the decision thresholds and driving agents’ best-responses, maintaining improvement incentives across all groups.
In conducting our analysis, we held the assumption of similar strategic behavior resources (cost and boost distributions) across groups. A main direction of future work includes exploring the impacts of varying these conditions among groups (e.g., availability of better strategic resources for the advantaged group), and the choice of interventions by the firm (e.g., when the firm makes the improvement action cheaper for a disadvantaged group group by providing free resources that help in improving skills/features), and make it more expensive for other groups. Similarly, this paper focused on studying agents’ strategic behavior when all agents (qualified and unqualified, and regardless of group membership) follow the same type of equilibrium best-responses as characterized in Proposition 2. Varying the resources available to different agents would also lead them to adopt different best-response types It would therefore be interesting to extend our analysis to scenarios where different groups adopt different types of equilibrium responses.
@@ A.1 Proofs for Section 4
Next, note that the set of (now) qualified agents with feature ${x}$ consists of the previously qualified agents with the same old feature ${x}$ (who have opted for $w=N$), the previously qualified or unqualified agents with feature $x-b$ who improved to feature $x$ (i.e., opted for $w=I$ and got a boost realization $b$), and previously qualified agents with feature $x-b$ who chose manipulation and reached feature ${x}$ (i.e., opted for $w=M$ and got a boost realization $b$). Thus,
$\hat{G}^{1}_{s}(x)=$ $\tfrac{\alpha_{s}}{\hat{\alpha}_{s}}\big{(}\mathbbm{1}(x\in\mathbb{N}^{1}_{s})G^{1}_{s}(x)+\int_{b}\mathbbm{1}(x-b\in\mathbb{M}^{1}_{s})G^{1}_{s}(x-b)\tau_{M,s}^{1}(b)\mathrm{d}b$
$+\int_{b}\mathbbm{1}(x-b\in\mathbb{I}^{1}_{s})G^{1}_{s}(x-b)\tau_{I,s}^{1}(b)\mathrm{d}b\big{)}+\tfrac{1-\alpha_{s}}{\hat{\alpha}_{s}}\int_{b}\mathbbm{1}(x-b\in\mathbb{I}^{0}_{s})G^{0}_{s}(x-b)\tau_{I,s}^{0}(b)\mathrm{d}b.$ (10)
We can re-write the expressions for the integrals using a change of variable $z:=x-b$ as follows:
$\int_{0}^{\infty}\mathbbm{1}(x-b\in\mathbb{M}^{1}_{s})G^{1}_{s}(x-b)\tau_{M,s}^{1}(b)\mathrm{d}b=\int_{x}^{-\infty}\mathbbm{1}(z\in\mathbb{M}^{1}_{s})G^{1}_{s}(z)\tau_{M,s}^{1}(x-z)\mathrm{d}(x-z)$
$=$ $\int_{-\infty}^{x}\mathbbm{1}(z\in\mathbb{M}^{1}_{s})G^{1}_{s}(z)\tau_{M,s}^{1}(x-z)\mathrm{d}z\int_{z\in\mathbb{M}^{1}_{s}}G^{1}_{s}(z)\tau_{M,s}^{1}(x-z)\mathrm{d}z~{}.$
@@
$\hat{G}^{0}_{s}(x)=$ $\tfrac{1-\alpha_{s}}{1-\hat{\alpha}_{s}}\big{(}\mathbbm{1}(x\in\mathbb{N}^{0}_{s})G^{0}_{s}(x)+\int_{b}\mathbbm{1}(x-b\in\mathbb{M}^{0}_{s})G^{0}_{s}(x-b)\tau_{M,s}^{0}(b)\mathrm{d}b\big{)}$ (11)

### v2

@@ 6.1 Effects of fairness interventions
We next proceed to studying how enforcing fairness at the policy level shapes agents’ strategic responses. As we show in this section, fairness constraints can have nontrivial and sometimes negative effects on improvement incentives across groups.
Specifically, assume the firm imposes a fairness constraint $\mathcal{C}$ when selecting the optimal threshold (as formalized in (3)), on the contrast between the unfair policies (characterized in Lemma 4 and 5) with the corresponding fair policies (characterized in Lemmas 6 and 7). Similar to Section 6, we do so when all agents opt for the same response type, and for Types 1 and 3 of agents’ best responses. Our comparison, outlined in the following Corollary, is a direct consequence of the proof of Proposition 8 under the stated assumptions.
@@
[…]
@@
In our context, Corollary 9 argues that in its attempt to attain fair outcomes, the firm decreases improvement incentives of the disadvantaged group $b$, while increasing the improvement incentives for the advantaged one; this can be viewed as a negative consequence of fairness interventions on shaping agents’ strategic responses. As we next elaborate through numerical experiments in Section 7, this negative by-product is more severe if the firm is non-strategic (i.e., unaware of agents’ strategic responses).
@@ 7 Numerical Experiments
In this section, we use numerical experiments to validate the analytical findings from Section 6. We focus on the affordable-improvement (Type 1) equilibrium, which features both manipulation and improvement; results for the costly-improvement (Type 3) case are deferred to Appendix C.3. Experiments are conducted on a semi-synthetic dataset based on real-world FICO credit data that includes four demographic groups: African American (AA), Hispanic (H), Caucasian (C), and Asian (A). To illustrate the main observations from Proposition 8, we first compare strategic vs. non-strategic policies in a two-group setting (AA vs. C) across varying target qualification rates $\alpha_{s}$. The full extension to all four groups is given in Appendix C.2.1. We then examine how fairness interventions reshape these incentive effects in settings with two demographic groups (C vs. AA), and subsequently extend the analysis to all demographic groups in Appendix C.2.2. Similar and additional experiments on a synthetic dataset are provided in Appendix D.
FICO data [Hardt et al., 2016].
We use the empirical FICO score distributions from prior work [Xie et al., 2024b, Zhang et al., 2022] to generate group-conditioned samples $(X,Y)\mid S=s$, without imposing any parametric structure. The dataset includes four demographic groups (AA, H, C, A) with qualification rates $\alpha_{s}\in(0.33849,0.56977,0.75972,0.80467)$, and all scores are normalized to $[0,1]$. The empirical CDFs, PDFs, and repayment likelihoods $P(Y=1\mid X,S)$ are shown in Appendix C.1 for reference.
Improvement vs. manipulation: empirical grounding.
[…]
Experimental setup for two demographic groups.
We compare policy behavior, without considering fairness, across two demographic groups: African American (AA) and Caucasian (C). For each group and target qualification rate $\alpha_{s}$, a single replicate proceeds as follows. We sample $n=1000$ agents such that approximately $\alpha_{s}n$ agents are drawn from the empirical distribution of scores among repayers ($y=1$) and the remaining agents are drawn from the empirical distribution among non-repayers ($y=0$). We vary $\alpha_{s}$ over $\{0.1,0.2,\dots,0.9\}$. For each $\alpha_{s}$ and each group, we run 50 independent replications and report averages across runs.
@@
Figure 2 shows that incorporating strategic responses reshapes behavior for both groups. We start the discussion on the unqualified agents’ behavior. Relative to a non-strategic firm, the strategic firm induces more genuine improvement (Figure 2(f)), with the effect being more pronounced for smaller values of $\alpha_{s}$. Simultaneously, it reduces overall manipulation (Figure 2(g)), although a slight increase is observed when $\alpha_{s}$ is very small (<0.25<0.25). This is because the firm has a considerable opportunity for inducing improvement in these settings. These effects appear in both AA and C groups and reflect a shift from gaming toward genuine improvement, which results in shifting the group’s qualification rates toward 1, as illustrated in Figure 2(c).
Although the non-strategic firm generates modest improvement, which is most noticeable around qualification rate of $0.25$ due to the high thresholds as appears in Figure 2(a), it rejects most unqualified agents who become qualified by improvement by missing out on their change due to being unaware of strategic behavior. As a result, downstream utility and improvement effects are weaker under non-strategic decision-making (Figures 2(b) and 2(f)).
(a) Optimal thresholds
(b) Optimal utility
(c) Post-strategic $\hat{\alpha}_{s}$
(d) $\bm{\Psi}^{1}_{s,(1)}$ (Manipulation)
(e) $\bm{\Phi}^{1}_{s,(1)}$ (Improvement)
(f) $\bm{\Phi}^{0}_{s,(1)}$ (Improvement)
(g) $\bm{\Psi}^{0}_{s,(1)}$ (Manipulation)
(h) Legend
Figure 2: Unfair policies: non-strategic vs. strategic policies’ comparison when varying pre-strategic $\alpha_{s}$.
For qualified agents, the effects depend on the group’s initial composition $\alpha_{s}$ (Figures 2(d) and 2(e)). When a group is majority qualified (high $\alpha_{s}$), the strategic firm induces both more improvement and manipulation; when it is majority unqualified (low $\alpha_{s}$), both improvement and manipulation decline as a result of the lower threshold (compared to non-strategic as seen in Figure 2(a)), thereby admits most qualified agents without additional effort. Note that the small corner divergences in the qualified-agents patterns relative to Proposition 8 arise because some of its technical assumptions (notably Assumption 3) are relaxed in these experiments. Importantly, both manipulation and improvement by such agents increase firm payoff. As a result, these deviations do not alter our main conclusions, which are driven by the behavior of unqualified agents and remain fully consistent with the theory. Consistent with this, Figure 2(b) shows that the strategic firm’s utility is significantly higher than that of the non-strategic firm.
A further implication is that the post-strategic qualification rate $\hat{\alpha}_{s}$ rises more for AA than for C when the initial composition is majority unqualified (Figure 2(c)). This is driven by the pre-strategic score distributions: most unqualified AA agents lie well below the threshold (e.g., many scores fall below $0.5$, which near the optimal strategic threshold as seen in Figure 2(a)), whereas many unqualified C agents are already near or above it ($0.6$) (See Appendix C.1). Consequently, a larger share of AA agents must undertake genuine improvement to be accepted, producing a larger upward shift in $\hat{\alpha}_{s}$ under strategic behavior.
(a) Strategic firm’s utility
(b) Non-strategic firm’s utility
Figure 3: Firm’s utility and optimal thresholds when imposing fairness constraints across Caucasian and African American groups.
Impacts of fairness interventions for strategic vs. non-strategic firms.
Figure 3 provides contour visualizations of the firm’s utility surface, together with the corresponding utilities under DP and EOP parity constraints, as the group-specific thresholds $\theta_{C}$ and $\theta_{AA}$ vary. The optimal thresholds for the unfair, DP-fair, and EOP-fair cases are also indicated in the figure. A key distinction is that the strategic firm evaluates fairness over post-strategic statistics, whereas the non-strategic firm evaluates fairness over pre-strategic statistics. Consequently, the fairness desiderata of a non-strategic firm are not satisfied ex-post, even though they appear satisfied under its own (incorrect) evaluation.
We first observe that, consistent with both prior literature and Corollary 9, fairness constraints induce both strategic and non-strategic firms to raise the threshold for the majority-qualified group and lower it for the majority-unqualified group. We also observe that the strategic firm (under both fair and unfair policies) selects higher thresholds than the non-strategic firm, in line with Proposition 8.
The more substantive distinction arises when comparing the type of fairness intervention. In particular, the gap between strategic and non-strategic firms is much larger under DP fairness than under EOP fairness. The DP-fair threshold chosen by a non-strategic firm (Figure 3(b)) is substantially lower for AA group, because the firm does not anticipate agents’ strategic responses. This leads to many agents in AA group being accepted by default, while others with even weaker features pass the lowered threshold through manipulation.
In contrast, a strategic firm lowers the AA group threshold under DP fairness only modestly (Figure 3(a)), recognizing that parity in selection rates can be achieved by jointly adjusting thresholds and inducing improvement responses. Consequently, fairness can be satisfied with far smaller changes to the decision rule, yielding a significantly smaller loss in utility once strategic behavior is accounted for.
@@
Impacts of fairness interventions on agents’ post-strategic qualification rates.
@@
To examine how strategic policies shape agents’ behavior under different fairness interventions (unfair, DP, and EOP), Figure 4 reports the post-strategic qualification rates $\hat{\alpha}_{s}$ for each group. As noted earlier, the strategic firm is consistently more effective at incentivizing improvement: both the C group and the AA group achieve higher $\hat{\alpha}_{s}$ under the strategic policy than under the non-strategic one (in Figure 4, the blue bars are always higher than the red bars). This effect is even more pronounced for AA group, reflecting its initially lower qualification rate.
More importantly, we observe that when a firm ignores strategic responses, the application of fairness interventions can have unintended negative consequences for the disadvantaged group (AA). Specifically, fairness constraints reduce the extent of improvement by AA group, as the red bars under fair policies fall below those under the unfair policy in Figure 4. The mechanism is straightforward: fairness constraints typically lower the decision threshold for AA group, which shifts behavior toward manipulation rather than genuine improvement. By contrast, a strategic firm internalizes agents’ responsiveness and uses fairness constraints together with threshold adjustment to continue incentivizing improvement, thereby generating higher post-strategic qualification rates for the disadvantaged group (AA).
Figure 4: Impact of fair vs. unfair and non-strategic vs. strategic policies on post-strategic $\hat{\alpha}_{s}$ across Caucasian and African American groups
@@ 8 Conclusion and Future Work
We proposed a Stackelberg framework for strategic classification, where a firm deploys a (fair) classifier, and agents strategically respond. Agents respond by either genuinely improving their qualification and observable features or by manipulating observable features to increase their chances of acceptance. We model heterogeneity in costs, and stochastic effectiveness of these actions yields three types of agents’ best responses based on cost–effectiveness relationships. We find that anticipating these strategic behaviors enables the firm to choose policies that both curb manipulation behavior (as also noted in prior work) and incentivize agents, particularly unqualified ones, to opt for improvement decisions. This shifts unqualified agents away from manipulation towards improvement, yet continues to allow some manipulation by qualified agents who would otherwise be rejected.
We further showed that imposing fairness constraints without anticipating agents’ strategic responses can inadvertently reduce improvement incentives for disadvantaged groups, since fairness may be achieved by accepting more individuals from this group. This can be viewed as a negative byproduct of fairness interventions. At the same time, such constraints can increase improvement incentives for advantaged groups as a result of a stricter threshold policy. In contrast, a strategic firm can satisfy fairness while largely maintaining incentives across all groups as it achieves fairness with minimal deviation from the unfair policy and by shaping agents’ best responses.
Our analysis assumes similar strategic behavior resources (cost and boost distributions) across groups, and focuses on settings in which all agents (qualified and unqualified) follow the same type of equilibrium best-responses, as characterized in Proposition 2. A natural direction for future work is to relax this assumption by allowing asymmetry, which would lead to heterogeneous response types across agents and groups, and yield a richer picture of how incentive design interacts with fairness in practice. An additional promising direction is to extend the model to allow endogenous costs that are shaped by the firm’s policy.
@@ A.1.2 Proof of Lemma 3
Next, note that the set of (now) qualified agents with feature x{x} consists of the previously qualified agents with the same old feature x{x} (who have opted for $w=N$), the previously qualified or unqualified agents with feature xbx-b who improved to feature $x$ (i.e., opted for $w=I$ and got a boost realization $b$), and previously qualified agents with feature xbx-b who chose manipulation and reached feature x{x} (i.e., opted for $w=M$ and got a boost realization $b$). Thus,
$\hat{G}^{1}_{s}(x)=$ $\tfrac{\alpha_{s}}{\hat{\alpha}_{s}}\big(\mathds{1}(x\in\mathbb{N}^{1}_{s})G^{1}_{s}(x)+\int_{b}\mathds{1}(x-b\in\mathbb{M}^{1}_{s})G^{1}_{s}(x-b)\tau_{M,s}^{1}(b)\mathrm{d}b$
$+\int_{b}\mathds{1}(x-b\in\mathbb{I}^{1}_{s})G^{1}_{s}(x-b)\tau_{I,s}^{1}(b)\mathrm{d}b\big)+\tfrac{1-\alpha_{s}}{\hat{\alpha}_{s}}\int_{b}\mathds{1}(x-b\in\mathbb{I}^{0}_{s})G^{0}_{s}(x-b)\tau_{I,s}^{0}(b)\mathrm{d}b.$ (10)
We can re-write the expressions for the integrals using a change of variable z:=xbz:=x-b as follows:
$\int_{0}^{\infty}\mathds{1}(x-b\in\mathbb{M}^{1}_{s})G^{1}_{s}(x-b)\tau_{M,s}^{1}(b)\mathrm{d}b=\int_{x}^{-\infty}\mathds{1}(z\in\mathbb{M}^{1}_{s})G^{1}_{s}(z)\tau_{M,s}^{1}(x-z)\mathrm{d}(x-z)$
$=$ $\int_{-\infty}^{x}\mathds{1}(z\in\mathbb{M}^{1}_{s})G^{1}_{s}(z)\tau_{M,s}^{1}(x-z)\mathrm{d}z\int_{z\in\mathbb{M}^{1}_{s}}G^{1}_{s}(z)\tau_{M,s}^{1}(x-z)\mathrm{d}z~.$
@@
$\hat{G}^{0}_{s}(x)=$ $\tfrac{1-\alpha_{s}}{1-\hat{\alpha}_{s}}\big(\mathds{1}(x\in\mathbb{N}^{0}_{s})G^{0}_{s}(x)+\int_{b}\mathds{1}(x-b\in\mathbb{M}^{0}_{s})G^{0}_{s}(x-b)\tau_{M,s}^{0}(b)\mathrm{d}b\big)$ (11)